# DEFT: Differentiable Branched Discrete Elastic Rods for Modeling Furcated DLOs in Real-Time

### v1

@@ V-A Hardware Setup
Figure 6: An illustration of experiment setup. Top: Data collection setup with a dual-arm robot. Bottom: An illustration of the BDLOs used to evaluate and compare the performance of DEFT against baseline methods. To increase dataset diversity, we vary the BDLOs’ stiffness, length, weights, number of vertices, and junction structures.
@@
TABLE I: A summary of each BDLO’s material properties and marker count used in the real-world experiment. Stiffness is estimated based on a relative scale.
@@ 4 Medium / 12 High / 2 High / 2 495 N/A
V-A1 Hardware Setup
Four distinct BDLOs were constructed to validate DEFT’s modeling accuracy as shown at the bottom of Figure 6. The material stiffness description of each BDLO is outlined in Table I. Note that we introduced stiff child branches attached to the parent branch in BDLO3 and BDLO4. These branches represent the rigid components (e.g., fixations) commonly used in BDLO construction [37]. These attachments help ensure secure mounting, preserve structural integrity, and manage loads in real‐world applications. Spherical markers were attached to each BDLO to facilitate dataset creation for both training and evaluation, with the number of vertices set equal to the number of MoCap markers. Because the markers are present during both training and evaluation, their impact on the BDLO dynamics is inherently accounted within all subsequent analyses. If one wishes to obtain a BDLO dataset without relying on real-world motion capture, a finite element method-based simulator could be employed. Note that the markers are only used to create training and evaluation datasets, as well as evaluate the performance of DEFT by providing the ground truth. However, for real-world applications, the markers are not needed. For dual manipulation, a Franka Emika Research 3 robot and a Kinova Gen3 robot were used. An illustration of the experimental setup is shown in Figure 6. We used an OptiTrack motion capture system to record ground-truth vertex locations. Note in our implementation of DEFT we set $\epsilon=0.1$ (Theorem 4) and $\kappa=0.02$ (Algorithm 2).
V-A2 Dataset Collection and Training
We collected dynamic trajectory data for selected BDLOs in the real world using a motion capture system operating at 100 Hz, capturing both slow and fast BDLO motions, as illustrated at the top of Figure 6. We split our data collection into two categories:
  • Dataset 1: each robot grasps one end of the BDLO. Data from this category were collected for all BDLOs.
  • Dataset 2: One robot grasped one end of the BDLO while the other robot grasped the BDLO at its midpoint. This dataset is collected for BDLOs 1 and 3.
Total collection times are provided in Table I. The dataset is split with $\sim$75for training and $\sim$25for evaluation. The model for each of the four BDLOs is trained separately using the trajectory data collected from our real-world experiments. During training, models are trained using the same 100 steps (i.e., 1s) of dynamic motion Each method’s performance is evaluated when predicting the state of the wire over a 500 step (i.e. 5s) prediction horizon using just the state of the wire at the initial time and the orientation and position of the two end effectors at each time step.
V-B Baseline Comparisons
Only a limited number of papers have investigated the dynamic behavior of BDLOs, and none provide a differentiable framework for learning real-world dynamics from real-world datasets. Moreover, no public code repositories are available for these approaches, making direct replication infeasible. We select two available implementations, Tree-LSTM [38] and GCN [33] as baseline prediction models for BDLOs. Tree-LSTM and GCN both capture structured relationships. Specifically, Tree-LSTM leverages hierarchical tree representations, while GCN excels in graph-based modeling. Because BDLOs can be described by nodes and edges, these architectures are naturally able to incorporate a BDLO’s dynamics.
V-C Modeling Results
Table II summarizes the average RMSE loss when comparing DEFT to baseline methods for two categories of dataset. Performance is evaluated over a 5 second prediction horizon for four different BDLOs. Note that DEFT outperforms all baselines across all four BDLOs. An illustration of the performance of DEFT, Tree-LSTM, and an ablation of DEFT, which is described in more detail in the next subsection, can be found in Figure 7. A video of related experiments can be found in the supplementary material.
TABLE II: Modeling Results Over a $5$s Horizon
Modeling Accuracy (RMSE, $10^{-2}$ m)
Dataset 1 Dataset 2
Method 1 2 3 4 1 3
GCN [33] 6.70 6.42 5.25 4.20 7.33 4.24
Tree-LSTM [38] 5.32 4.50 2.46 2.71 7.08 4.55
DEFT 1.87 2.82 1.51 1.41 1.49 1.80
V-D Compuational Experiment
TABLE III: One Step Inference Time with Dataset 1
Time ($10^{-2}$ s)
Method 1 2 3 4
GCN [33] 0.20 0.19 0.13 0.13
Tree-LSTM [38] 1.81 2.44 1.79 1.89
DEFT 0.76 0.70 0.91 0.93
@@ V-D Compuational Experiment
This section evaluates the computational efficiency of DEFT when compared to baseline methods. It also examines how employing the analytical gradient derived from Theorem 1 to compute (10) and the parallel programming approach detailed in Section IV-F, impacts overall computational time. All experiments in this section were conducted in Python on an Ubuntu 20.04 equipped with an AMD Ryzen PRO 5995WX CPU, 256GB of RAM, and 128 cores.
V-D1 Computational Speed Comparison
We compare computational speed for one step inference. Table III indicates that GCN has the fastest speed. DEFT is the second fastest and is able to generate single step predictions at a 100Hz frequency.
V-D2 Computational Speed with Analytical Gradient
We compare the computational speed of using analytical gradients when compared to using numerical gradients. As summarized in Table IV, computing (10) with the analytical gradient is approximately twice as fast as using numerical gradients.
TABLE IV: Compuational Time for (10) with Dataset 1
@@ V-D2 Computational Speed with Analytical Gradient
V-D3 Computational Speed with Parallel Programming
We compare the computational performance of using parallel programming approach described in Section IV-F to a naive (sequential) implementation, in which each branch is simulated one after the other. In this experiment, we use the material properties of BDLO1 and gradually increase the number of children branches from 1 to 8. As shown in Figure 9, the naive implementation’s computation time grows more rapidly than that of the parallel approach. However, a certain increase in computational time is still expected with parallelization. As illustrated in Figure 3, constraints at each junction between the parent branch and its child branches prevent a fully batch-based operation, thereby resulting in an increase in the computational time.
TABLE V: Ablation Study with Dataset 1
Modeling Accuracy (RSME, $10^{-2}$ m)$\downarrow$
Dataset 1 Dataset 2
Method 1 2 3 4 1 2
W/O Constraints.2, 3 and 4 $7.66\times 10^{8}$ $6.10\times 10^{8}$ $6.56\times 10^{6}$ $1.77\times 10^{8}$ $3.81\times 10^{7}$ $1.23\times 10^{8}$
W/O Constraints.3 and 4 $1.36\times 10^{2}$ $1.81\times 10^{2}$ $1.51\times 10^{2}$ $7.64\times 10^{1}$ $8.72\times 10^{1}$ $3.38\times 10^{2}$
W/O Constraint.4 3.95 3.49 2.03 2.15 4.01 4.88
W/O Learning + System ID 2.46 3.53 1.85 1.77 2.03 2.21
W/O System ID 2.24 3.00 1.67 1.50 1.66 1.95
W/O Residual Learning 1.93 3.19 1.70 1.62 1.86 1.89
Full Model 1.87 2.82 1.51 1.41 1.49 1.80
V-E Ablation Study
The results presented in Table V demonstrate the contribution of each DEFT component. Notably, excluding Theorems 2, 3, and 4 causes a significant increase in prediction loss. This outcome is primarily driven by the absence of Theorem 2, which leads to simulation instability. Although residual learning is used to mitigate errors, directly modeling stiff behaviors, such as the inextensibility of DLOs, remains challenging, making the learning process highly sensitive to inputs and hampering effective gradient propagation. A key challenge in using neural networks to learn position-based constraints is that constraints such as inextensibility confine the system to a low-dimensional manifold within a high-dimensional space. Without these constraints, it becomes difficult for the network alone to keep the system on that manifold. Over multiple open-loop predictions, even minor errors can accumulate and push the system state off the manifold, causing reduce of residual learning’s effect. Similarly, removing Theorem 3 allows the children branch to detach from parent branch, and therefore fall freely under gravity, leading to a substantial increase in loss. Moreover, excluding Theorem 4 reduces accuracy, demonstrating the importance of introducing rigidity at junctions to allow dynamic propagation across branches.
Despite these challenges, auto-tuning the material properties provides the most substantial improvement over the baseline, illustrating DEFT’s ability to leverage real-world data for parameter identification. Residual learning also yields a non-trivial performance gain. By contrast, the model that do not learn from real-world data underperforms those that do, illustrating the importance of differentiability in this framework.
V-F Planning Results
This section demonstrates how these BDLO models can be used for planning tasks, including shape matching and thread insertion. We compare three approaches: DEFT, which is our proposed method; DEFT without the constraint described in Theorem 4, an ablation approach that removes rigidity at junctions and thus limits dynamic propagation across branches; Tree-LSTM, the baseline model with performance closest to DEFT. In these experiments, we employ ARMOUR [39], a receding-horizon trajectory planner and tracking controller, to manipulate BDLOs. ARMOUR relies on each model’s predicted BDLO states to guide the BDLO to a desired final configuration. Further details on integrating ARMOUR with BDLO simulators are provided in Appendix -E. The results are summarized in Table VI, and Figure 8 shows one of the real-world results. A video of these experiments can be found in the supplementary material.
TABLE VI: Real-World Trajectory Planning Results
Successful Rate
Shape Matching Thread Insertion
Tree-LSTM [38] 2/40 0/35
DEFT w/o Theorem 4 27/40 16/35
DEFT 35/40 29/35
V-F1 Shape Matching
In this experiment a goal configuration for the wire is specified and a random initial condition for the wire is chosen. A trial is considered successful if the Euclidean distance between each ground-truth vertex and its corresponding target vertex after planning is less than 0.05 m. If the task is not completed within 30 seconds, the trial is marked as a failure. As summarized in Table  VI (left column), DEFT achieves a higher success rate than the other two models. Note that these experiments were all conducted in the real-world which further demonstrates the practical applicability of DEFT.
V-F2 Thread Insertion
In this experiment, the planning algorithm must insert the end of the BDLO into a hole of radius $2.5$cm which is placed in arbitrary location. The process consists of two stages—shape matching and thread insertion—which together require two goal configurations for the wire. Unlike the previous experiment, in which the robot grasps the branch end, the robot’s end effector clamps the third vertex from the branch’s insertion point. The trial is deemed successful if the BDLO branch successfully passes through the torus while also matching the user-specified wire shape successfully. As shown in Table  VI (right column), DEFT again achieves a higher success rate than the other two models. Note that these experiments were all conducted in the real-world which further demonstrates the practical applicability of DEFT.
VI Limitations
While DEFT shows promising results in modeling and control of BDLOs, several limitations remain that could be addressed in future work. First, DEFT currently lacks a contact model for accurately predicting interactions between deformable components and their surroundings. Incorporating a contact modeling framework, such as that proposed in [40], would enable more realistic simulations and more robust planning in environments where collisions and frictional forces play a significant role. Second, for the insertion tasks considered in this work, we assume that the location and geometry of the target hole are known in advance. In unstructured or partially observed environments, this assumption may not hold. Incorporating 3D scene understanding methods, such as Gaussian Splatting [41], could help autonomously detect or refine the location of target features and thereby increasing the system’s applicability to real-world scenarios. Lastly, this work does not account for the orientation of the gripper during BDLO manipulation, which can significantly affect deformation behavior. As discussed in [16], gripper orientation can influence the kinematics of deformable objects. Future extensions of DEFT should therefore incorporate a more comprehensive treatment of gripper orientation to improve accuracy and expand the range of manipulable tasks.
VII Conclusion
This paper introduces DEFT, a method that embeds residual learning within a novel differentiable branched Deformable Linear Object (BDLO) simulator. DEFT enables accurate modeling and real-time prediction of a BDLO’s dynamic behavior over long time horizons under dynamic manipulation. To demonstrate DEFT’s efficacy, we conduct a comprehensive set of experiments evaluating its accuracy, computational speed, impacts of its constribution and utility for manipulation tasks. Compared to the state of the art, DEFT achieves higher accuracy while remaining sample-efficient and without compromising computational speed. This paper also illustrates how to integrate DEFT with a planning and control framework for 3D shape matching and thread insertion of a BDLO in real-world scenarios. DEFT successfully completes more complex, real-world manipulation tasks than other tested baselines.
Acknowledgement

### v5

@@ C.1 Hardware Setup
Figure 9: An illustration of experiment setup. Left: Data collection setup with a dual-arm robot. Right: An illustration of the BDLOs used to evaluate and compare the performance of DEFT against baseline methods. To increase dataset diversity, we vary the BDLOs’ stiffness, length, weights, number of vertices, and junction structures.
@@
Table 4: A summary of each BDLO’s material properties and marker count used in the real-world experiment. Stiffness is estimated based on a relative scale.
@@ 4 Medium / 12 High / 2 High / 2 495 N/A
C.2 Training Setup
Let $\mathbf{U}_{1:T-1}$ denote the set of inputs applied from time $t=1$ to $t=T-1$, and let $\mathbf{X}_{1:T}=\{\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{T}\}\in\mathbb{R}^{T\times n\times 3}$ denote the corresponding ground-truth trajectory of the BDLO from $t=1$ to $t=T$. Given known initial conditions $\mathbf{X}_{1}$, $\mathbf{V}_{1}$, and inputs $\mathbf{U}_{1:T-1}$, Algorithm 1 can be recursively applied to predict the trajectory $\hat{\mathbf{X}}_{2:T}$. Let $\bm{\phi}$ denote the parameters of the GNN. The training objective is formulated as the following optimization problem:
$\underset{\bm{\alpha},\bm{\phi}}{\min}\sum_{t=1}^{T-1}\|\mathbf{X}_{t+1}-\hat{\mathbf{X}}_{t+1}\|_{2}$ (56)
Leveraging DEFT’s differentiability, $T$ can be selected as greater than 2 to better capture long-term dynamics. This multi-step training approach results in improved prediction accuracy compared to single-step training, as demonstrated in Table 5.
Table 5: Ablation Study with Dataset 1
Modeling Accuracy (RSME, $10^{-2}$ m)$\downarrow$
Dataset 1
Method           1          2
Single-step Training           2.22          3.41
Multi-step Training           1.87           2.82
C.3 Computational Speed
@@
This section evaluates the computational efficiency of DEFT when compared to baseline methods. It also examines how employing the analytical gradient derived from Theorem 2 to compute (3) and the parallel programming approach detailed in Section 4.5, impacts overall computational time. All experiments in this section were conducted in Python on an Ubuntu 20.04 equipped with an AMD Ryzen PRO 5995WX CPU, 256GB of RAM, and 128 cores.
C.3.1 Computational Speed with Analytical Gradient
Table 6: Compuational Time for (3) with Dataset 1
@@ C.3.1 Computational Speed with Analytical Gradient
We compare the computational speed of using analytical gradients when compared to using numerical gradients. As summarized in Table 6, computing (3) with the analytical gradient is approximately twice as fast as using numerical gradients.
C.3.2 Computational Speed with Parallel Programming
We compare the computational performance of using parallel programming approach described in Section 4.5 to a naive (sequential) implementation, in which each branch is simulated one after the other. In this experiment, we use the material properties of BDLO1 and gradually increase the number of children branches from 1 to 8. As shown in Figure 10, the naive implementation’s computation time grows more rapidly than that of the parallel approach. However, a certain increase in computational time is still expected with parallelization. As illustrated in Figure 6, constraints at each junction between the parent branch and its child branches prevent a fully batch-based operation, thereby resulting in an increase in the computational time.
C.4 Ablation Study on Residual Learning Architecture
To further demonstrate the significance of using GNN for residual learning, we replace it with the fully connected MLP described in Appendix B.3 and test on BDLO 2 and 3 from Dataset 1. We use BDLO 2 and 3 from dataset 1. As shown in Table 7, residual learning using a naive MLP is running slightly more efficient but provides limited improvements when compared to the proposed GNN approach. This illustrates the importance of the GNN because even modest improvements are important for prediction accuracy in long-term horizon predictions, where small errors can be significantly amplified during extended open-loop predictions.
Table 7: Ablation Study Comparing GNN and MLP Residual Learning Approaches
Accuracy ($10^{-2}$ m) Comp. Time ($10^{-2}$ s)
Method BDLO 2 BDLO 3 BDLO 2 BDLO 3
Replace GNN with MLP 3.11 1.64 0.67 0.72
DEFT 2.82 1.51 0.70 0.76
C.5 Modeling and Planning Visualization
Additional visulization of modeling and planning results can be found in Figure 11 and Figure 12 respectively.
@@ C.5 Modeling and Planning Visualization
Appendix D ARMOUR
To address the shape matching task, we employ ARMOUR [33], an optimization-based framework for motion planning and control. The objective is to guide a robot arm so that it manipulates a BDLO from an initial configuration to a specified target configuration. ARMOUR accomplishes this through a receding-horizon approach: at each iteration, it solves an optimization problem to determine the robot’s trajectory. ARMOUR parameterizes the robot’s joint trajectories as polynomials, with polynomial coefficients serving as the decision variables. One end of the rope is rigidly attached to the robot’s end effector, whose pose is determined by forward kinematics. By substituting the end effector’s final pose into DEFT, we obtain a prediction of the rope’s configuration at the conclusion of the motion. The optimization’s cost function then minimizes the Euclidean distance between this predicted BDLO configuration from DEFT and the target configuration. To ensure the resultant motion is physically feasible, ARMOUR incorporates constraints on the robot’s joint positions, velocities, and torques. Once the optimization is solved, the resulting trajectory is tracked by ARMOUR’s controller. This planning–execution process is repeated until the rope achieves the desired configuration or a maximum iteration threshold is exceeded (in which case the task is deemed a failure). Additional details on ARMOUR’s trajectory parameterization and its closed-loop controller can be found in [33, Section IX].